OpenStudy (anonymous):

find the difference 4y/xy - 2x/xy^3

OpenStudy (anonymous):

your least common multiple of \(xy\) and \(xy^3\) is \(xy^3\) so you need to multiply top and bottom of first fraction by \(y^2\) and then you can add straight across

OpenStudy (anonymous):

4y/xy-2x/xy^3 =4/x-2/y^3 =(4y^3-2x)/xy^3